OpenStudy (anonymous):

-x^2+5x =3 use the quadratic formula to solve

OpenStudy (shamil98):

move everything to one side and use the quadratic formula then..

OpenStudy (jack1):

-x^2 + 5x =3 -x^2 + 5x -3 =0

OpenStudy (jack1):

in this example -x^2 + 5x -3 =0 ax^2 + bx +c =0 so a = -1 b = 5 c = -3